OpenStudy (anonymous):

Vocabulary from Classical Roots, Book C, Exercise 7c is tricky... Please help!!! (: 1.Bulbs planted during the fall _____________ remain until spring. 3. Greek temples such as Parthenon were built as ______________ for deities 5.Lady Macbeth , who walks through the castle washing her hands in her sleep , is perhaps the most famous ____________________ in literature. 6. The student is fortunate who is as comfortable in the _______________ of math and science as in that of the humanities. 7.The armies of Alexander the Great established _________________ over the entire Middle East.

OpenStudy (anonymous):

The choices are… ablution, deluge, divest, domain, domestic, domicile, domineer, dominion, dormant, investiture, somnambulate, somnolent, travesty, vested, vestment,

OpenStudy (anonymous):

Select the correct reasons for statements 1 and 3 to complete the proof. Prove: 4x + 8 • x • 3 = 28x 1. 4x + 8 • x • 3 = 4x + 8 • 3 • x 1. ? 2. 4x + 8 • x • 3 = 4x + 24x 2. Multiplication 3. 4x + 8 • x • 3 = 28x 3. ? A. Associative Property of Addition, Addition B. Commutative Property of Addition, Multiplication C. Associative Property of Multiplication, Addition D. Commutative Property of Multiplication, Addition